OpenStudy (anonymous):

Which is a factor of 16x^3 + 250 ? Choose one answer. a. 2x - 5 b. 4x^2 + 10x + 25 c. 4x^2 + 20x + 25 d. 2x + 5

sam (.sam.):

What do you think?

OpenStudy (anonymous):

b

sam (.sam.):

You can factor 2 out of that f(x)=2(8x^3+125) Then trial by error if you prefer, f(-5/2)=0 f(-5/2)=8(-5/2)^3+125=0 So -5/2 is a factor of f(x)

sam (.sam.):

Then you get 2x+5

sam (.sam.):

x=-5/2 2x=-5 2x+5=0

sam (.sam.):

You there?